### v1

# Hypergeometric Distributions and joint families of elliptic curves Brian Grove and Hasan Saad

Abstract.
Recently, the first author as well as the second author with Ono, Pujahari, and Saikia determined the limiting distribution of values of certain finite field and hypergeometric functions. These hypergeometric values are related to Frobenius traces of elliptic curves and their limiting distribution is determined using connections to the theory of modular forms and harmonic Maass forms. Here we determine the limiting distribution of values of some hypergeometric functions which are sums of traces of Frobenius for a pair of elliptic curves. To obtain this result, we generalize Michel’s work on Sato–Tate laws for families of elliptic curves to the setting of pairs of families, and we show that a generic pair admits an independent Sato–Tate distribution as the finite field grows. To this end, we use various results from the theory of étale cohomology, Deligne’s work on the Weil conjectures, and the work of Katz on monodromy groups. In the cases previously studied using modular methods, we elucidate the connection between the modular forms that appear and the machinery of étale cohomology.
Key words and phrases:
Gaussian hypergeometric functions, Sato–Tate type distributions, Elliptic curves
2020 Mathematics Subject Classification:
11G20, 11T24, 33E50
1. Introduction and Statement of Results
In the 1980’s, Greene [Gre84, Gre87] defined hypergeometric functions over finite fields as an analogue of the classical hypergeometric functions. Using the analogy between character sum expansions and power series expansions, he proved various special evaluations and transformation formulas that closely mirror their classical counterparts. These functions have played central roles in the study of combinatorial supercongruences [Ahl01, AO00a, KC17, McC12a, MOS20, OSZ18], elliptic curves [BK13, Fus10, Len11a, McC13, Ono98, TY18], surfaces [AOP02, DKS+20, Ono98], Dwork hypersurfaces [BRS16, McC17, Goo17], hyperelliptic curves [BKS14, BSM15], Calabi–Yau threefolds [LTYZ21, AO00b], the theory of hypergeometric motives and Galois representations [RRV22, HT20, AGLT11, AGLT16, FLR+22], and the theory of modular forms [FOP04, Fus13, FM16, Len11b, MP15, Pap06, HLLT18].
The limiting distributions of these finite field hypergeometric functions have been studied in multiple works [OSS23, Saa23, OPSS24, Gro24]. To make this precise, we first recall the definition of these finite field hypergeometric functions. Let be a prime, a fixed primitive th root of unity, and let be a generator of the character group of Furthermore, let and be two collections of rational numbers and let be the greatest common denominator of 111We say that the hypergeometric datum is of length In this notation, if and then we have [McC12b] the finite field hypergeometric function222The function is a variant of the hypergeometric functions originally defined by Greene.
where
is the Gauss sum corresponding to Furthermore, we define In this notation, the aforementioned works determine the limiting distribution of as and varies over for certain hypergeometric data of lengths two and three.
Here we determine the limiting behavior of for certain hypergeometric data of length four for which the condition can be removed (see Theorem 1.3 of [BCM15]). The following theorem gives the limiting moments.
Theorem 1.1.
Let and If then we have
where is the -th Catalan number.
Remark 1.2.
The Catalan numbers are ubiquitous in mathematics ([Sta15]). In relation to our results, we have that these Catalan numbers
arise (see (5.3) of [FKRS12]) as the even moments of the traces of the Lie group More precisely, if then we have
where the integral is taken with respect to the Haar measure. Equivalently, is (see Lemma 4.3) the multiplicity of the trivial representation in the direct sum decomposition of the -th tensor power representation of
As a corollary, we obtain the limiting distribution of the values as varies over and in terms of the Meijer -functions (see Definition 7.2).
Corollary 1.3.
Let and If then we have
Example.
For the prime the histogram illustrates the case of Corollary 1.3, i.e. the near match with the limiting distribution.
histogram for and
Geometrically, the arithmetic of the 333In what follows, we denote by the hypergeometric functions in Theorem 1.1 with understood from context. functions in Theorem 1.1 is intimately related to the arithmetic of certain families of elliptic curves (see Table 1). To make this precise, if is a prime, and then we have that (see Lemma 2.3)
where and where
In order to prove Theorem 1.1, we require a different set of tools than those used in [OSS23, Saa23, OPSS24, Gro24]. To make this precise, we shortly recall the method of proof in those works. To this end, recall that if then the th Chebyshev polynomial of the second kind is defined as the unique polynomial such that
More explicitly, we have that
In this notation, if and then the authors of [OSS23, Saa23, OPSS24, Gro24] use either the theory of harmonic Maass forms or the theory of Hecke traces for triangle groups in [HLLT18] to express the twisted th moment
as a Fourier coefficient of a holomorphic cusp form. By Deligne’s bound on the Fourier coefficients of cusp forms and induction, we have that if then the limiting -th moment is given by
An application of the method of moments then gives the required distributions. In contrast to the methods described above, the moments of the functions we consider, or equivalently the mixed moments of and cannot be described in terms of modular forms, which necessitates most of the work in this paper.
Here we use (1.1) in order to write the moments of as weighted combinations of mixed moments of the traces and In order to apply the method of moments, we determine the asymptotics of the mixed moments for generic pairs of families of elliptic curves. To make this precise, for let and consider the two families of elliptic curves defined by
The family can be put in a Weierstrass form (see III.1 of [Sil09])
where In this notation, we let and be the discriminant and -invariant of respectively. We say that is a generic pair if the following hold:
and are nonconstant.
There exists an such that for the reductions of and mod have either good or multiplicative reduction at each
There exists an such that for and the reductions of and mod are not isogenous over where
is the quadratic twist of by
Finally, for and a prime write In this notation, we compute the limiting mixed moments of and
Theorem 1.4.
If is a generic pair and then we have
where is the th Catalan number.
Using these moments, we determine the limiting joint distribution of the pair
Corollary 1.5.
If is a generic pair, and then we have
Remark 1.6.
The conditions of a generic pair of elliptic curves which is required for Theorem 1.4 and Corollary 1.5 can be relaxed (see Proposition 5.1) to the condition that
for any rank -adic system where is the projection map For example (see Lemma 3.10 and proof of Lemma 6.4), if has additive potentially multiplicative reduction for some while has good reduction, this condition would be satisfied. In general, if one allows additive reduction, then the situation is more delicate and one needs to take into account twists by nonconstant rational functions.
We remark that the methods used to prove Theorem 1.4 and Corollary 1.5 immediately offer a different - more geometric - proof of the limiting moments of the hypergeometric functions of length studied in [OSS23, OPSS24, Gro24]. Moreover, it also gives the case which is not accessible through the modular methods used in these works.
Theorem 1.7.
If and then we have
Finally, with very slight modifications to our method, we offer an alternative proof of the limiting moments of the hypergeometric function of length studied in [OSS23, Saa23].
Theorem 1.8.
If then we have
Remark 1.9.
We give the following remarks regarding Theorems 1.4, 1.7, 1.8, and Corollary 1.5.
We remark that Theorem 1.4 and Corollary 1.5 are well-known to the experts (for example, see Chapters 9 and 10 of [KS99]). The work in this paper illustrates how the geometric results obtained by Deligne [Del80], Katz [Kat88, Kat90a], and others serve to obtain arithmetic-statistical results in number theory, as well as to explain the appearance of Catalan numbers in [OSS23, Saa23, OPSS24, Gro24]. Moreover, our methods, when restricted to the cases studied with modular methods, yield an explicit description of the “mysterious” cohomology groups that appear within this paper, elucidating the connection between the modular methods and the abstract machinery of étale cohomology in the situations where such a connection exists (through a universal elliptic curve over a modular curve; for example, see Sections 3 and 4 of [HLLT18]).
We remark that Theorem 1.7 follows immediately from the work of Michel [Mic95]. We add it here for completeness and to illustrate the connection between the geometric methods used here and the modular methods used in [OSS23, OPSS24, Gro24].
The methods used to prove Theorem 1.8 explicitly bound the difference between the actual and limiting moments which leads to an explicit bound on the difference between the limiting and actual distributions of the functions à la [Saa23]. It is important to note that this yields no extra power-savings in the error compared to modular methods.
Geometrically, the arithmetic of the functions in Theorem 1.8 is related to twisted symmetric squares of the Clausen family of elliptic curves (see Lemma 2.2).
The moments in Theorem 1.8 arise (see [PV04]) as the even moments of the traces of the Lie group More precisely, if is even, then we have
Our proof of Theorem 1.4 closely mirrors the work of Michel on the equidistribution of Frobenius traces for a generic family of elliptic curves [Mic95, Proposition 1.1]. First, we describe and as traces of Frobenius on two -adic sheaves and lisse on an open subset In this notation, the -th mixed moments are then described as sums of traces of Frobenius on the sheaves Under the conditions on a generic pair, we decompose these sheaves as a direct sum of irreducible representations of the so-called geometric monodromy group. The multiplicities of the trivial representation in these direct sum decompositions are either or products of Catalan numbers. This allows us to explicitly compute the contribution of the trivial parts to the moments. The contribution of the non-trivial parts is then shown to be small using the Grothendieck–Lefschetz trace formula and Deligne’s [Del80] results on the purity of cohomology groups in his seminal proof of the generalized Weil conjectures. The proof of Theorem 1.7 follows from the intermediate propositions that lead up to proving Theorem 1.4, whereas Theorem 1.8 follows from the same methods with slight modifications.
This paper is organized as follows. In Section 2, we recall important facts on the relations between the functions mentioned above and families of elliptic curves. In Section 3, we recall necessary background from the theory of étale cohomology and important facts on the first cohomology of elliptic curves. In Section 4, we recall facts from representation theory. In Section 5, we apply the results from étale cohomology and representation theory to compute the mixed moments of families of elliptic curves under a certain “independence” condition of their first cohomologies. In Section 6, we describe this “independence” condition in terms of isogenies of families of elliptic curves viewed as elliptic curves over In Section 7, we recall the method of moments and derive the necessary distributions for Corollaries 1.3 and 1.5. In Section 8, we conclude with the proofs of Theorems 1.1 through 1.8 and Corollaries 1.3 and 1.5.
Acknowledgements
The authors would like to thank Ling Long and Fang-Ting Tu for the numerous discussions and advice which highly contributed to the quality of this paper.
The authors would also like to thank Bill Hoffman and Ken Ono for their multiple comments, discussions, and helpful remarks. The authors also thank Wanlin Li who pointed out an error in an earlier draft. The first author was partially supported by a research assistantship from the Louisiana State University (LSU) Department of Mathematics.
2. The functions and the arithmetic of elliptic curves
Here we recall important facts about the relations between functions and the arithmetic of one-parameter families of elliptic curves.
We first recall the relationship between the hypergeometric functions in Theorem 1.7 and elliptic curves. If and is prime, define the families of elliptic curves by
Table 1.
In this notation, the functions give the finite field point counts on these elliptic curves.
Finally, we relate the functions in Theorem 1.1 to the elliptic curves
Lemma 2.3.
If is prime, and then we have when is not a square in Moreover, if then we have
Proof.
If is not a square, the fact that follows immediately from Theorem 2.2 and Lemma 2.6 of [MT24].
To prove (2.1) when note that Corollary 2.3 and Lemma 2.6 of [MT24] imply that
In order to prove Theorems 1.1 through 1.8, we describe the traces of an elliptic curve as traces of Frobenius morphisms on an -adic étale sheaf. This perspective allows us to use the resuls and heavy machinery of étale cohomology which we assume the reader is familiar with. In Subsection 3.1, for the reader’s convenience, we recall the necessary background from the theory of étale cohomology. In Subsection 3.2 we recall some background on the sheaves which encode finite-field point counts on families of elliptic curves. Those sheaves “are” essentially the rational -adic Tate modules when the family is viewed as an elliptic curve defined over the function field (see Section 6).
Remark 3.1.
For background on the applications of étale cohomology in analytic number theory, the reader is referred to [FKMS19], [FKM14], [FKM15], [Kat90b]. For a more geometric perspective, the reader is referred to [Mil80].
3.1. Étale sheaves and monodromy groups
Here we recall results from the theory of étale cohomology. More precisely, the results we recall allow us to view powers of Frobenius traces at different fibers simultaneously, and to then rewrite the sums of those powers in terms of the traces on the so-called cohomology group. For example, in [OSS23, OPSS24, Gro24], certain weighted sums of moments of Frobenius traces are expressed explicitly as Fourier coefficients of certain cusp forms. The first cohomology groups that appear here should be thought of as (see Remark 5.4) the geometric generalization of these forms when one does not have an underlying modular setting. We closely follow [FKMS19]. In what follows in this subsection, let be a prime, the field of rational functions over and let be a prime.
Recall that if then an -adic sheaf that is lisse on “is” (more accurately, see Section 1 of Chapter V of [Mil80]) a continuous finite-dimensional Galois representation
where is a finite-dimensional -vector space of dimension and where is unramified at points To make this precise, recall that for every we have a decomposition group an inertia group a residue field and an exact sequence
In this notation, we say that is unramified at if acts trivially on Furthermore, for each we denote by the Frobenius morphism
and by the inverse of In this notation, for each we have a lift of to an -class of Furthermore, we denote by the geometric Frobenius on namely, the inverse of the arithmetic Frobenius
In order to prove Theorems 1.1 and 1.4, we will describe the moments of the values as sums of the form where is an appropriately chosen étale sheaf. To determine the asymptotics of these moments, we express the sums over as sums of traces of acting on cohomology groups associated to the sheaf
To make this precise, if denote by the base-change of to If is an -adic sheaf that is lisse on we have the associated cohomology groups with compact support These groups are finite-dimensional -representations of In this notation, the Grothendieck–Lefschetz trace formula expresses the sums of traces of Frobenius over a sheaf as an alternating sum of traces of Frobenius on the cohomology groups.
Theorem 3.2(Th. 4.1 of [FKMS19], Th. 13.4 of [Mil80]).
If is lisse on then
In order to estimate the sums we require bounds on both the eigenvalues of Frobenius and on the dimension of the cohomology groups. To this end, let be an -adic sheaf lisse on with We say that is punctually pure of weight if for each the eigenvalues of are such that for every embedding In this notation, the fundamental theorem of Deligne (see III of [Del80]) bounds the eigenvalues of Frobenius on the middle cohomology.
If is an -adic sheaf lisse on that is punctually pure of weight and is an eigenvalue of then for all embeddings
It turns out that the extremal cohomology groups and are easy to express in terms of the representation of the geometric Galois group. To make this precise, let We then have an induced exact sequence
where with the compositum of and We say that is geometrically irreducible (resp. geometrically constant) if is irreducible (resp. trivial) as a representation of Furthermore, if is an -adic sheaf lisse on and then up to taking a finite extension of we denote by the “half”-Tate twist (see Remark 3.11 of [FKMS19]) of by In this notation, we have the following description of and
Lemma 3.4(Section 4.1 of [FKMS19], Remark 2.4 of [Mil80]).
If is an -adic sheaf lisse on then we have that
and
where is the largest quotient module of on which acts trivially. In particular, if is geometrically irreducible and nonconstant, then
In our application of the Grothendieck–Lefschetz trace formula, the sheaves we consider are irreducible. In this case, Lemma 3.4 implies that we only require the dimension of the middle cohomology The Grothendieck–Ogg–Shafarevich formula gives the alternating sum of the dimensions of the cohomology groups in terms of local data attached to the sheaf.
If is an -adic sheaf lisse on then the Euler characteristic of is given by
where is the Swan conductor of at In particular, if is irreducible and has at most tame ramification on we have that
In order to prove our distribution results, we will show that the sheaves we consider are irreducible as representations of To this end, we require the notion of geometric monodromy group. More precisely, if is a lisse sheaf on associated to the representation then the geometric monodromy group is the Zariski closure (for the definition and properties of Zariski closure, see Section 7.1 of [Kow14]) of in The sheaves we consider for the computation of the mixed moments arise as tensor products (in the representation-theoretic sense) of two sheaves. A special case of the Goursat–Kolchin–Ribet criterion allows us to compute the geometric monodromy groups of these tensor products.
Theorem 3.6(Th. 13.4 of [FKMS19], 1.8.2 of [Kat90a]).
Let and be two generically rank- geometrically irreducible -adic sheaves such that the following hold:
and are punctually pure of weight and self-dual (as representations).
The geometric monodromy groups of and are
For any rank one sheaf -adic sheaf we have that is not geometrically isomorphic to that is, not isomorphic as representations of .
Then, the geometric monodromy group of the sheaf is
Finally, note that in Lemma 2.2, the relation between the function and the symmetric square of involves a Legendre character This meshes well with the framework of étale cohomology as the Kummer sheaf has traces equal to values of multiplicative characters.
Lemma 3.7.
[FKMS19, Section 3.4 and Example 4.2.2]
Let be a prime a character on and Then there exists a rank- -adic sheaf , pure of weight , lisse on and tamely ramified at and such that
for all
3.2. Étale sheaves arising from families of elliptic curves
In his proof of equidistribution of traces of Frobenius for a generic family of elliptic curves [Mic95, Section 3], Michel considers certain étale sheaves which contain information about those traces. Here, we recall the construction and properties of these sheaves for the reader’s convenience.
Let be a family of elliptic curves defined by the affine variety
equipped with the projection map
In this notation, we define to be the Tate twist of the first higher direct image with proper support of the constant sheaf on (for the definition of higher direct images with proper support, see Chapter VI Section 3 of [Mil80]).
Let be the discriminant of and let Furthermore, let such that the leading coefficient of is a unit in Then, we have that and that is lisse on
Now denote by and the base-change of and to Then, is an -adic sheaf, lisse on pure of weight and for we have
where In order to compute the moments and mixed moments, we require the geometric monodromy group, the self-duality, and the Euler characteristic of
Lemma 3.9.
There exists such that for all primes the following are true.
If the -invariant is not constant as a function of then the geometric monodromy group of is given by
is tamely ramified at the points of
The sheaf is self-dual.
Sketch of proof.
Parts (1) and (2) are Lemma 3.1 of [Mic95]. To see part (3), note that the smooth base change theorem (see Chapter VI, Corollary 4.2 of [Mil80]) implies that is isomorphic as a Galois representation to the where is viewed as an elliptic curve over The self-duality then follows by Poincaré duality (see Chapter VI, Corollary 11.2 of [Mil80]) which in this case is nothing but the Weil pairing.
Finally, in order to apply Theorem 3.6 in the case of generic pairs of families of elliptic curves, it is useful to compute the action of on where is a point of bad reduction of
Lemma 3.10(p. 135 of [Mic95], Section 4 of [Del80]).
Let such that the fiber has multiplicative reduction. Then, in some basis, the representation restricted to is given by
where is the canonical surjection from to
4. Background from Representation Theory
Here we recall facts from representation theory which allow for computation of moments using the representations described in Section 3. Aside from the technical details, this section is essentially a generalization of some of the tools used in [OSS23, OPSS24, Gro24]. More precisely, the results in this section combined with Lemma 3.9 elucidate why the Chebyshev polynomial sums
are precisely the sums with small asymptotic magnitude. Furthermore, this section elucidates why the Catalan numbers are the moments and transforms the induction arguments on Chebyshev polynomials into direct explicit formulas.
In order to compute the moments of traces of Frobenius we make use of the tensor product representations of
Definition 4.1.
If is a group, and is a finite-dimensional representation of the tensor product representation is given by
where
From the definition, it is obvious that In order to apply the results from Section 3 to where is the standard representation of we decompose into a direct sum of irreducible representations. To this end, recall that if and is a finite-dimensional representation, then we have the -th symmetric power representation
where is the quotient of by the natural action of the symmetric group and where acts naturally on this quotient space. In this notation, we have the following important facts when is the standard representation of
If is the standard representation of on a vector space over a field of characteristic and then is an irreducible representation of and for all we have
where is the -th Chebyshev polynomial of the second kind.
In order to compute the moments in Theorems 1.1 and 1.7, we decompose the tensor product representations into a direct sum of symmetric power representations and a trivial representation. The aim of this is to isolate the irreducible nontrivial components, which by Lemma 3.4, Theorem 3.2, and Theorem 3.3 have small contributions to the moments. The following well-known lemma (for example, see 2.2. of [LZ10]) provides the required direct sum decomposition.
Lemma 4.3.
If is the standard representation of and then we have
where
In particular, when is even, the multiplicity of the trivial representation is given by the Catalan number
In [OSS23, OPSS24, Gro24], the authors compute the moments by an induction argument on the contributions of the Chebyshev polynomials of the traces of Frobenius. Lemma 4.3 serves as a direct expression of the -th moments in terms of the various Chebyshev polynomials that appear in those works. For example, if then we have
This, for example, makes it clear why the statement that
for all proved in the aforementioned works, implies that the -th moment is given by the Catalan number
In our work, we only make use of the multiplicity of the trivial representation in The full decomposition is noted only for completeness.
In the setting of Theorem 1.8, Lemma 2.2 implies that if is even, then the -th moments of the functions are normalized even moments of In order to compute these moments, we decompose the tensor product representation of into a direct sum of symmetric power representations and a trivial representation. For the even moments, we require the following lemma which gives the multiplicity of the trivial representation.
Lemma 4.5.
If is the standard representation of and , then the multiplicity of the trivial representation in the direct sum decomposition of is given by
Proof.
If then this is trivial. Now suppose
Using Lemma 4.3, we have that
where is the trivial representation. Taking -th tensor powers on both sides and comparing the multiplicities of the trivial representation, we have that
where is the multiplicity of the trivial representation in the direct sum decomposition of Therefore, if we define
then, by induction on our claim is equivalent to showing that
Expanding the definition of this is equivalent to showing that
By switching the order of summation, we write the left-hand side as
However, we have that the inner sum is given by
which by the binomial theorem is clearly equal to
Putting all this together, we have our claim.
For the odd moments, we require the moments of twisted by the trace of a Kummer sheaf. To compute those moments, we require the following elementary lemma (see Exercise 2.2.14 of [Kow14] for example) on the irreducibility of tensor products with one-dimensional representations.
Lemma 4.6.
If is a group, a finite-dimensional vector space over a field an irreducible representation of and a one-dimensional representation of then is irreducible.
In order to use the geometric monodromy computation of Deligne in Lemma 3.9, we connect the irreducibility of with
Recall that the geometric mondromy group is the Zariski closure of in The following well-known lemma (for example, see Exercise 7.1.6 of [Kow14]) relates the irreducibility of a representation
to the irreducibility of the (standard) representation of its Zariski closure in
Lemma 4.7.
Let be a group and be a finite-dimensional linear representation of We have that is irreducible if and only if the standard representation of is irreducible.
Proof.
Here we write the proof for the reader’s convenience.
If is irreducible, it follows immediately that acts irreducibly. Now, suppose the standard representation of is irreducible, and let be a subspace of such that
Let be the subset of elements such that
It is clear that is a Zariski closed algebraic subgroup of which contains This implies that Since acts irreducibly, or proving our claim.
The following technical lemma plays an essential role in deducing irreducible decomposition of tensor products of sheaves from the geometric monodromy computation in Theorem 3.6.
Lemma 4.8.
Let be a group, be finite-dimensional representations of and let be Zariski-continuous irreducible representations of the Zariski closures respectively. If
then is an irreducible representation of
Proof.
The assumption
implies that
Since the pair is Zariski-continuous, we have that
where is the external tensor product (see p. 64 of [Kow14]) of and Since and are irreducible, we have that (see Proposition 2.3.23 of [Kow14]) is irreducible, and therefore, the standard representation of is irreducible. By Lemma 4.7, we conclude that is an irreducible representation of
Finally, in order to verify the conditions of Theorem 3.6 in our setting by applying Lemma 3.10, we compute the space of invariants after twisting by a rank- -adic sheaf.
Lemma 4.9.
Let be a group and be a rank- representation of such that under some basis, for every we have
where for all In this notation, if is a nontrivial rank- representation of then the space of invariants of is
Proof.
Let be the basis in the statement of the lemma, and let be an element of If the subspace generated by is invariant under the action of a direct computation shows that
5. Moments of Frobenius Traces
Here we assemble the results from Sections 3 and 4 to compute the moments and mixed-moments that are required for Theorems 1.1 through 1.8.
To make this precise, for let be the two affine varieties over given by
with Furthermore denote by the projection map
and define For a prime we denote by the base-change of to We further assume that there the -invariants of and are nonconstant, and that there exists such that for all and all rank- -adic sheaves on we have
Finally, write In this notation, we have the following preliminary version of Theorem 1.4.
Proposition 5.1.
If then we have
Note that the case computes the moments of one-parameter families of elliptic curves with nonconstant -invariant (without need of choice for by the same proof), recovering the result of Michel [Mic95, Section 3], of which this is a generalization.
In order to prove this proposition, we first compute the multiplicity of the trivial representation in the -adic sheaves To make this precise, let be the subset of on which both and are lisse (see Lemma 3.8) and for each prime denote by and the base-changes of to and respectively. In this notation, we have the following computation.
Lemma 5.2.
If and then the multiplicity of the trivial representation in the direct-sum decomposition of as a representation of is given by
Proof.
Denote by the representation
As representations of if we have the decompositions
where the representations are irreducible pairwise nonisomorphic nontrivial representations of and where is the trivial representation. Part (1) of Lemma 3.9 states that the Zariski closure of Therefore, by Lemma 4.7, we have that the are precisely the symmetric power representations and are the multiplicities in Lemma 4.3.
Now, note that by Theorem 3.6, we have that the Zariski closure of is given by Therefore, Lemma 4.8 implies that for we have that
is an irreducible representation of It is also clear that and are irreducible.
Therefore, the multiplicity of the trivial representation in the decomposition of is equal to and then Lemma 4.3 shows our claim.
The next lemma computes the asymptotics of trace sums of the nontrivial irreducible representations arising in the direct sum decomposition of More explicitly, the proof of Lemma 5.2 shows that these representations are of the form where with appropriate multiplicities. Therefore, the following lemma proves that if and are not both zero, then we have
Lemma 5.3.
If is a prime and then we have
where runs over the nontrivial irreducible representations arising in the direct sum decomposition of
Proof.
If is an irreducible representation as in the statement, it is clear that viewed as an -adic sheaf, is lisse on Therefore, Theorem 3.2 gives us that
In the proof of Lemma 5.3, we show that as we have
by identifying the sum with the trace of Frobenius on the cohomology group where
Now, for let and let be the family defined in Section 2. In this case, we have an explicit description of the Frobenius traces on this “mysterious” cohomology group in terms of the traces of Hecke operators on spaces of cusp frorms. For example, if and is even, then (see Section 4.2 of [Gro24]) the trace of the Frobenius map is given by
where is the trace of the -th Hecke operator on the space of cusp forms of weight over and where is an explicit expression in terms of The same remark applies to odd and to the elliptic curves The advantage of the modular method is that one understands the trace of Frobenius on the cohomology very explicitly. On the other hand, the advantage of the machinery of étale cohomology is that one can derive results on distributions in a more general setting where there is no underlying modular framework.
where is the multiplicity of the trivial representation in Combining this with Lemma 5.2, we obtain our claim.
In order to obtain the necessary moments for Theorem 1.8, we require the following modification of Proposition 5.1.
Proposition 5.5.
If then we have
Proof.
If is even, then we have that
Using Lemma 5.2, we decompose the right-hand sum as
where is the sheaf associated to the Clausen curves as in Section 3.2 and where ranges over with and Using exactly the same argument as in the proof of Lemma 5.3, we obtain our claim.
If is odd, note that
where is the Kummer sheaf associated to the Legendre character Decomposing as in the case for even moments and using the orthogonality of characters, we have that
Now, note that these irreducible representations are symmetric powers of By Lemma 4.6, we have that is irreducible. Moreover, it is clear that this representation is never trivial. Therefore, the same argument as in the proof of Lemma 5.3 shows our claim.
Remark 5.6.
Just as in Proposition 5.1, the cohomology groups that show up in the proof of Proposition 5.5 have explicit traces using modular forms. For more details, see Proposition 2.2 and the proof of Proposition 6.3 in [Saa23].
Finally, to compute the moments in Theorem 1.1 in terms of the mixed moments in Theorem 1.4, we require the following combinatorial lemma.
Lemma 5.7.
If then we have
Proof.
If by definition of the Catalan numbers, we have
Expanding the right-hand side, we then have that
We rewrite this expression as
An application of the Chu–Vandermonde identity (see Example 1.17 of [Jac13]) then gives that
which proves our claim.
6. Tate Modules and Isogenies
Here we discuss the connection between the (untwisted) sheaves discussed in Section 3.2 and -adic Tate modules, and then express the non-isomorphism condition required for Proposition 5.1 in terms of isogenies.
First, if is a field, is an elliptic curve, and is prime, recall that the rational -adic Tate module is defined by
and that admits a natural action of the underlying Galois group (see III.7 of [Sil09]). It turns out that the sheaves we consider “are” nothing but duals of those rational Tate modules when is viewed as an elliptic curve over
where is viewed as an elliptic curve over and where the isomorphism is an isomorphism of representations of the Galois group
Proof.
This is Theorem 15.1 of [Mil86] and Corollary 4.2 of [Mil80].
The following result of Paršin plays an essential role in showing that a generic pair of elliptic curves satisfies the conditions required to use Proposition 5.1.
Let be a prime and let be a function field of transcendence degree over where If and are elliptic curves over and then and are isogenous over
Now, let be as in Section 5. The following proposition describes the “independence” condition of a generic pair in terms of isogenies of elliptic curves.
Proposition 6.3.
Let be a generic pair. There exists such that if and is a rank- -adic sheaf on then we have
In order to prove Proposition 6.3, we require the following two lemmas.
Lemma 6.4.
Assume and have either good or multiplicative reduction everywhere. There exists such that if is a prime and is a rank- -adic sheaf on with
then is geometrically constant.
Proof.
First, by Lemma 3.10 and our assumption of good or multiplicative reduction, it is clear that and must have the same points of ramification. By Lemma 4.9, we have that must be lisse on Furthermore, by Lemma 3.9, there exists such that if , then and are tamely ramified at and therefore, is tamely ramified at Therefore, is geometrically constant (see Example 5.2 (f) of Chapter I of [Mil80]).
Lemma 6.5.
Assume that and have either good or multiplicative reduction everywhere. There exists such that if is a prime and is a rank- -adic sheaf on with
then we have
where is the projection map on the quadratic twist and
Proof.
If is geometrically constant, then it follows by Proposition 2.8.2 of [Kow14] and (3.2) that for we have
where is a character on By observing that the traces of Frobenius and are -valued for all embeddings of and by the purity of weight of the sheaves we have that is either a trivial or a quadratic character.
Now, take to be a nonsquare in and consider By comparing the traces of Frobenius of and the quasi-orthogonality relations of irreducible sheaves (see Theorem 5.2 of [FKMS19]) imply that, up to enlarging we have for all primes
This immediately follows from Lemmas 6.1, 6.4, 6.5, and Theorem 6.2.
7. Some Distributions
To obtain Corollaries 1.3 and 1.5, we combine Theorems 1.1 and 1.4 with the method of moments. In this section, we compute the required probability density functions associated to a given sequence of moments or mixed moments.
For the mixed moments, the following lemma gives the probability density functions.
Lemma 7.1.
If is prime and consider maps
Suppose that for each we have that
If then we have
Sketch of Proof.
The method of moments (see Exercises 30.5 and 30.6 of [Bil12]) states that in order to compute the limiting distribution, under very general settings, it suffices to show that the limiting moments coincide with the moments of the claimed distribution. In our case, this is precisely the statement that if then we have that
which is straightforward.
We now determine the probability density function for the moments arising in Theorem 1.1. To this end, we recall the definition of the Meijer -function (see Appendix of [EMOT54]).
Definition 7.2.
Let be integers and suppose are complex numbers such that for all and Furthermore, suppose where
In this notation if and the Meijer -function is given as
where denotes Euler’s Gamma function, and where is a loop beginning and ending at that encircles all poles of for exactly once in the positive direction and does not encircle any poles of for
In order to compute integrals involving Meijer -functions, we require the following key transformation and integral identities.
Lemma 7.3.
The following are true.
If then we have
Under general conditions (see 20.5 of [EMOT54]), if then we have
We will also make use of the following elementary lemma which expresses products of Catalan numbers in terms of ratios of Gamma functions.
Lemma 7.4.
If then we have
Proof.
Using the property that and that we have that
where Therefore, we have that
and our claim follows immediately.
Finally, we compute the probability density function associated to the moments that appear in Theorem 1.1.
Lemma 7.5.
If is prime, consider maps
Suppose that for each we have that
If then we have
Proof.
By the method of moments (see Theorems 30.5 and 30.6 of [Bil12]), it suffices to show that for we have
If is odd, this is clear since the integrand is odd. If the integrand is even and we rewrite the integral as
Using (1) of Lemma 7.3 and the change-of-variables we have
Computing the integral using (2) of Lemma 7.3, we obtain
has multiplicative reduction at whereas has good reduction at Therefore, Proposition 5.1 applied to combined with Lemma 5.7 gives the claimed moments.
This follows from Lemma 2.2 and Proposition 5.5.

### v2

@@
-Hypergeometric Distributions and joint families of elliptic curves Brian Grove and Hasan Saad
+Hypergeometric Distributions and joint families of elliptic curves
 Abstract.
@@
 Key words and phrases:
-Gaussian hypergeometric functions, Sato–Tate type distributions, Elliptic curves
-2020 Mathematics Subject Classification:
-11G20, 11T24, 33E50
-1. Introduction and Statement of Results
+Gaussian hypergeometric functions, Sato–Tate type distributions, Elliptic curves2020 Mathematics Subject Classification:
+11G20, 11T24, 33E501. Introduction and Statement of Results
 In the 1980’s, Greene [Gre84, Gre87] defined hypergeometric functions over finite fields as an analogue of the classical hypergeometric functions.
@@
 To make this precise, we first recall the definition of these finite field hypergeometric functions.
-Let be a prime, a fixed primitive th root of unity, and let be a generator of the character group of Furthermore, let and be two collections of rational numbers and let be the greatest common denominator of 111We say that the hypergeometric datum is of length In this notation, if and then we have [McC12b] the finite field hypergeometric function222The function is a variant of the hypergeometric functions originally defined by Greene.
+Let be a prime, a fixed primitive th root of unity, and let be a generator of the character group of Furthermore, let and be two collections of rational numbers and let be the least common multiple of the denominators of 111We say that the hypergeometric datum is of length In this notation, if and then we have [McC12b] the finite field hypergeometric function222The function is a variant of the hypergeometric functions originally defined by Greene.
 where
 is the Gauss sum corresponding to Furthermore, we define In this notation, the aforementioned works determine the limiting distribution of as and varies over for certain hypergeometric data of lengths two and three.
-Here we determine the limiting behavior of for certain hypergeometric data of length four for which the condition can be removed (see Theorem 1.3 of [BCM15]).
+Here we determine the limiting behavior of for certain hypergeometric data of length four.
+Remark 1.1.
+For the hypergeometric data we consider, the condition can be weakened to (see Theorem 1.3 of [BCM15]).
+Therefore, since we are concerned with the asymptotic behavior as in what follows, we have no restrictions on
 The following theorem gives the limiting moments.
@@
 Remark 1.3.
-The Catalan numbers are ubiquitous in mathematics ([Sta15]). In relation to our results, we have that these Catalan numbers
-arise (see (5.3) of [FKRS12]) as the even moments of the traces of the Lie group More precisely, if then we have
+The Catalan numbers are ubiquitous in mathematics ([Sta15]).
+In relation to our results, we have that these Catalan numbers arise (see (5.3) of [FKRS12]) as the even moments of the traces of the Lie group More precisely, if then we have
 where the integral is taken with respect to the Haar measure.
@@
 For the prime the histogram illustrates the case of Corollary 1.4, i.e. the near match with the limiting distribution.
-histogram for and
 Geometrically, the arithmetic of the 333In what follows, we denote by the hypergeometric functions in Theorem 1.2 with understood from context.
@@
 Remark 1.10.
-We give the following remarks regarding Theorems 1.4, 1.7, 1.8, and Corollary 1.5.
 We remark that Theorem 1.5 and Corollary 1.6 are well-known to the experts (for example, see Chapters 9 and 10 of [KS99]).
@@
 Moreover, our methods, when restricted to the cases studied with modular methods, yield an explicit description of the “mysterious” cohomology groups that appear within this paper, elucidating the connection between the modular methods and the abstract machinery of étale cohomology in the situations where such a connection exists (through a universal elliptic curve over a modular curve; for example, see Sections 3 and 4 of [HLLT18]).
-We remark that Theorem 1.7 follows immediately from the work of Michel [Mic95].
-We add it here for completeness and to illustrate the connection between the geometric methods used here and the modular methods used in [OSS23, OPSS24, Gro24].
+- (2)
 The methods used to prove Theorem 1.9 explicitly bound the difference between the actual and limiting moments which leads to an explicit bound on the difference between the limiting and actual distributions of the functions à la [Saa23].
 It is important to note that this yields no extra power-savings in the error compared to modular methods.
-Geometrically, the arithmetic of the functions in Theorem 1.8 is related to twisted symmetric squares of the Clausen family of elliptic curves (see Lemma 2.2).
-The moments in Theorem 1.8 arise (see [PV04]) as the even moments of the traces of the Lie group More precisely, if is even, then we have
+- (4)
+- (5)
 Our proof of Theorem 1.5 closely mirrors the work of Michel on the equidistribution of Frobenius traces for a generic family of elliptic curves [Mic95, Proposition 1.1].
@@
 In Section 5, we apply the results from étale cohomology and representation theory to compute the mixed moments of families of elliptic curves under a certain “independence” condition of their first cohomologies.
-In Section 6, we describe this “independence” condition in terms of isogenies of families of elliptic curves viewed as elliptic curves over In Section 7, we recall the method of moments and derive the necessary distributions for Corollaries 1.3 and 1.5.
-In Section 8, we conclude with the proofs of Theorems 1.1 through 1.8 and Corollaries 1.3 and 1.5.
+In Section 6, we describe this “independence” condition in terms of isogenies of families of elliptic curves viewed as elliptic curves over In Section 7, we recall the method of moments and derive the necessary distributions for Corollaries 1.4 and 1.6.
+In Section 8, we conclude with the proofs of Theorems 1.2 through 1.9 and Corollaries 1.4 and 1.6.
 Acknowledgements
@@
 The authors also thank Wanlin Li who pointed out an error in an earlier draft.
+Finally, the authors thank the anonymous referee whose comments improved the quality of the exposition.
 The first author was partially supported by a research assistantship from the Louisiana State University (LSU) Department of Mathematics.
@@
 We first recall the relationship between the hypergeometric functions in Theorem 1.8 and elliptic curves. If and is prime, define the families of elliptic curves by
-Table 1.
 In this notation, the functions give the finite field point counts on these elliptic curves.
+For the hypergeometric datum the function is related to the finite field point counts on the Clausen elliptic curves
+Lemma 2.2 (Th. 5 of [Ono98]).
+If is prime and then we have
+where is the Legendre character on and where
 Finally, we relate the functions in Theorem 1.2 to the elliptic curves
@@
 If is prime, and then we have when is not a square in Moreover, if then we have
-Proof.
-If is not a square, the fact that follows immediately from Theorem 2.2 and Lemma 2.6 of [MT24].
-To prove (2.1) when note that Corollary 2.3 and Lemma 2.6 of [MT24] imply that
-In order to prove Theorems 1.1 through 1.8, we describe the traces of an elliptic curve as traces of Frobenius morphisms on an -adic étale sheaf.
+3. Background from Étale Cohomology
+In order to prove Theorems 1.2 through 1.9, we describe the traces of an elliptic curve as traces of Frobenius morphisms on an -adic étale sheaf.
 This perspective allows us to use the resuls and heavy machinery of étale cohomology which we assume the reader is familiar with.
@@
 Remark 3.1.
-For background on the applications of étale cohomology in analytic number theory, the reader is referred to [FKMS19], [FKM14], [FKM15], [Kat90b].
-For a more geometric perspective, the reader is referred to [Mil80].
 3.1. Étale sheaves and monodromy groups
@@
 and by the inverse of In this notation, for each we have a lift of to an -class of Furthermore, we denote by the geometric Frobenius on namely, the inverse of the arithmetic Frobenius
-In order to prove Theorems 1.1 and 1.4, we will describe the moments of the values as sums of the form where is an appropriately chosen étale sheaf.
+In order to prove Theorems 1.2 and 1.5, we will describe the moments of the values as sums of the form where is an appropriately chosen étale sheaf.
 To determine the asymptotics of these moments, we express the sums over as sums of traces of acting on cohomology groups associated to the sheaf
 To make this precise, if denote by the base-change of to If is an -adic sheaf that is lisse on we have the associated cohomology groups with compact support These groups are finite-dimensional -representations of In this notation, the Grothendieck–Lefschetz trace formula expresses the sums of traces of Frobenius over a sheaf as an alternating sum of traces of Frobenius on the cohomology groups.
-Theorem 3.2(Th. 4.1 of [FKMS19], Th. 13.4 of [Mil80]).
-If is lisse on then
 In order to estimate the sums we require bounds on both the eigenvalues of Frobenius and on the dimension of the cohomology groups.
 To this end, let be an -adic sheaf lisse on with We say that is punctually pure of weight if for each the eigenvalues of are such that for every embedding In this notation, the fundamental theorem of Deligne (see III of [Del80]) bounds the eigenvalues of Frobenius on the middle cohomology.
+Theorem 3.3 (Th. 4.6 of [FKMS19]).
 If is an -adic sheaf lisse on that is punctually pure of weight and is an eigenvalue of then for all embeddings
@@
 In this case, Lemma 3.4 implies that we only require the dimension of the middle cohomology The Grothendieck–Ogg–Shafarevich formula gives the alternating sum of the dimensions of the cohomology groups in terms of local data attached to the sheaf.
+Theorem 3.5 (Th. 4.2. of [FKMS19]).
 If is an -adic sheaf lisse on then the Euler characteristic of is given by
@@
 Lemma 3.7.
-[FKMS19, Section 3.4 and Example 4.2.2]
-Let be a prime a character on and Then there exists a rank- -adic sheaf , pure of weight , lisse on and tamely ramified at and such that
+[FKMS19, Section 3.4 and Example 4.2.2] Let be a prime a character on and Then there exists a rank- -adic sheaf , pure of weight , lisse on and tamely ramified at and such that
 for all
@@
 In this notation, we define to be the Tate twist of the first higher direct image with proper support of the constant sheaf on (for the definition of higher direct images with proper support, see Chapter VI Section 3 of [Mil80]).
+We first recall some elementary properties of
+Lemma 3.8 (Section 3 of [Mic95]).
 Let be the discriminant of and let Furthermore, let such that the leading coefficient of is a unit in Then, we have that and that is lisse on
@@
 To see part (3), note that the smooth base change theorem (see Chapter VI, Corollary 4.2 of [Mil80]) implies that is isomorphic as a Galois representation to the where is viewed as an elliptic curve over The self-duality then follows by Poincaré duality (see Chapter VI, Corollary 11.2 of [Mil80]) which in this case is nothing but the Weil pairing.
+∎
 Finally, in order to apply Theorem 3.6 in the case of generic pairs of families of elliptic curves, it is useful to compute the action of on where is a point of bad reduction of
-Lemma 3.10(p. 135 of [Mic95], Section 4 of [Del80]).
-Let such that the fiber has multiplicative reduction. Then, in some basis, the representation restricted to is given by
-where is the canonical surjection from to
 4. Background from Representation Theory
@@
 Aside from the technical details, this section is essentially a generalization of some of the tools used in [OSS23, OPSS24, Gro24].
-More precisely, the results in this section combined with Lemma 3.9 elucidate why the Chebyshev polynomial sums
-are precisely the sums with small asymptotic magnitude.
+More precisely, the results in this section combined with Lemma 3.9 elucidate why the Chebyshev polynomial sums are precisely the sums with small asymptotic magnitude.
 Furthermore, this section elucidates why the Catalan numbers are the moments and transforms the induction arguments on Chebyshev polynomials into direct explicit formulas.
@@
 In this notation, we have the following important facts when is the standard representation of
+Lemma 4.2 (Theorems 2.6.1 and 5.6.3 of [Kow14]).
 If is the standard representation of on a vector space over a field of characteristic and then is an irreducible representation of and for all we have
 where is the -th Chebyshev polynomial of the second kind.
-In order to compute the moments in Theorems 1.1 and 1.7, we decompose the tensor product representations into a direct sum of symmetric power representations and a trivial representation.
+In order to compute the moments in Theorems 1.2 and 1.8, we decompose the tensor product representations into a direct sum of symmetric power representations and a trivial representation.
 The aim of this is to isolate the irreducible nontrivial components, which by Lemma 3.4, Theorem 3.2, and Theorem 3.3 have small contributions to the moments.
@@
 In particular, when is even, the multiplicity of the trivial representation is given by the Catalan number
+Remark 4.4.
+We give the following two remarks on Lemma 4.3.
 In [OSS23, OPSS24, Gro24], the authors compute the moments by an induction argument on the contributions of the Chebyshev polynomials of the traces of Frobenius.
@@
 Proof.
-If then this is trivial. Now suppose
-Using Lemma 4.3, we have that
+If then this is trivial. Now suppose Using Lemma 4.3, we have that
 where is the trivial representation. Taking -th tensor powers on both sides and comparing the multiplicities of the trivial representation, we have that
@@
 which by the binomial theorem is clearly equal to
-Putting all this together, we have our claim.
+Putting all this together, we have our claim. ∎
 For the odd moments, we require the moments of twisted by the trace of a Kummer sheaf.
@@
 If is a group, a finite-dimensional vector space over a field an irreducible representation of and a one-dimensional representation of then is irreducible.
-In order to use the geometric monodromy computation of Deligne in Lemma 3.9, we connect the irreducibility of with
-Recall that the geometric mondromy group is the Zariski closure of in The following well-known lemma (for example, see Exercise 7.1.6 of [Kow14]) relates the irreducibility of a representation
+In order to use the geometric monodromy computation of Deligne in Lemma 3.9, we connect the irreducibility of with Recall that the geometric mondromy group is the Zariski closure of in The following well-known lemma (for example, see Exercise 7.1.6 of [Kow14]) relates the irreducibility of a representation
 to the irreducibility of the (standard) representation of its Zariski closure in
@@
 Here we write the proof for the reader’s convenience.
-If is irreducible, it follows immediately that acts irreducibly. Now, suppose the standard representation of is irreducible, and let be a subspace of such that
+If is irreducible, it follows immediately that acts irreducibly.
+Now, suppose the standard representation of is irreducible, and let be a subspace of such that
 Let be the subset of elements such that
-It is clear that is a Zariski closed algebraic subgroup of which contains This implies that Since acts irreducibly, or proving our claim.
+It is clear that is a Zariski closed algebraic subgroup of which contains This implies that Since acts irreducibly, or proving our claim. ∎
 The following technical lemma plays an essential role in deducing irreducible decomposition of tensor products of sheaves from the geometric monodromy computation in Theorem 3.6.
@@
 Since the pair is Zariski-continuous, we have that
+Putting (4.1) and (4.2) together, we have that
 where is the external tensor product (see p.
@@
 Proof.
-Let be the basis in the statement of the lemma, and let be an element of If the subspace generated by is invariant under the action of a direct computation shows that
+Let be the basis in the statement of the lemma, and let be an element of If the subspace generated by is invariant under the action of a direct computation shows that ∎
 5. Moments of Frobenius Traces
-Here we assemble the results from Sections 3 and 4 to compute the moments and mixed-moments that are required for Theorems 1.1 through 1.8.
+Here we assemble the results from Sections 3 and 4 to compute the moments and mixed-moments that are required for Theorems 1.2 through 1.9.
 To make this precise, for let be the two affine varieties over given by
@@
 Part (1) of Lemma 3.9 states that the Zariski closure of Therefore, by Lemma 4.7, we have that the are precisely the symmetric power representations and are the multiplicities in Lemma 4.3.
-Now, note that by Theorem 3.6, we have that the Zariski closure of is given by Therefore, Lemma 4.8 implies that for we have that
-is an irreducible representation of It is also clear that and are irreducible.
+Now, note that by Theorem 3.6, we have that the Zariski closure of is given by Therefore, Lemma 4.8 implies that for we have that is an irreducible representation of It is also clear that and are irreducible.
 Therefore, the multiplicity of the trivial representation in the decomposition of is equal to and then Lemma 4.3 shows our claim.
+∎
 The next lemma computes the asymptotics of trace sums of the nontrivial irreducible representations arising in the direct sum decomposition of More explicitly, the proof of Lemma 5.2 shows that these representations are of the form where with appropriate multiplicities.
@@
 Proof.
-If is an irreducible representation as in the statement, it is clear that viewed as an -adic sheaf, is lisse on Therefore, Theorem 3.2 gives us that
+Remark 5.4.
 In the proof of Lemma 5.3, we show that as we have
-by identifying the sum with the trace of Frobenius on the cohomology group where
-Now, for let and let be the family defined in Section 2.
+by identifying the sum with the trace of Frobenius on the cohomology group where Now, for let and let be the family defined in Section 2.
 In this case, we have an explicit description of the Frobenius traces on this “mysterious” cohomology group in terms of the traces of Hecke operators on spaces of cusp frorms.
@@
 On the other hand, the advantage of the machinery of étale cohomology is that one can derive results on distributions in a more general setting where there is no underlying modular framework.
-where is the multiplicity of the trivial representation in Combining this with Lemma 5.2, we obtain our claim.
+Now we prove Proposition 5.1.
+Proof of Proposition 5.1.
 In order to obtain the necessary moments for Theorem 1.9, we require the following modification of Proposition 5.1.
@@
 Therefore, the same argument as in the proof of Lemma 5.3 shows our claim.
+∎
 Remark 5.6.
-Just as in Proposition 5.1, the cohomology groups that show up in the proof of Proposition 5.5 have explicit traces using modular forms.
-For more details, see Proposition 2.2 and the proof of Proposition 6.3 in [Saa23].
 Finally, to compute the moments in Theorem 1.2 in terms of the mixed moments in Theorem 1.5, we require the following combinatorial lemma.
@@
 An application of the Chu–Vandermonde identity (see Example 1.17 of [Jac13]) then gives that
-which proves our claim.
+which proves our claim. ∎
 6. Tate Modules and Isogenies
@@
 It turns out that the sheaves we consider “are” nothing but duals of those rational Tate modules when is viewed as an elliptic curve over
+Lemma 6.1.
+If and are as in Section 3.2, then
 where is viewed as an elliptic curve over and where the isomorphism is an isomorphism of representations of the Galois group
-Proof.
-This is Theorem 15.1 of [Mil86] and Corollary 4.2 of [Mil80].
 The following result of Paršin plays an essential role in showing that a generic pair of elliptic curves satisfies the conditions required to use Proposition 5.1.
+Theorem 6.2 (Corollary 1 of Theorem 2 of [Par72]).
 Let be a prime and let be a function field of transcendence degree over where If and are elliptic curves over and then and are isogenous over
@@
 By Lemma 4.9, we have that must be lisse on Furthermore, by Lemma 3.9, there exists such that if , then and are tamely ramified at and therefore, is tamely ramified at Therefore, is geometrically constant (see Example 5.2 (f) of Chapter I of [Mil80]).
+∎
 Lemma 6.5.
@@
 where is a character on By observing that the traces of Frobenius and are -valued for all embeddings of and by the purity of weight of the sheaves we have that is either a trivial or a quadratic character.
-Now, take to be a nonsquare in and consider By comparing the traces of Frobenius of and the quasi-orthogonality relations of irreducible sheaves (see Theorem 5.2 of [FKMS19]) imply that, up to enlarging we have for all primes
-This immediately follows from Lemmas 6.1, 6.4, 6.5, and Theorem 6.2.
+Now, take to be a nonsquare in and consider By comparing the traces of Frobenius of and the quasi-orthogonality relations of irreducible sheaves (see Theorem 5.2 of [FKMS19]) imply that, up to enlarging we have for all primes ∎
 7. Some Distributions
-To obtain Corollaries 1.3 and 1.5, we combine Theorems 1.1 and 1.4 with the method of moments.
+To obtain Corollaries 1.4 and 1.6, we combine Theorems 1.2 and 1.5 with the method of moments.
 In this section, we compute the required probability density functions associated to a given sequence of moments or mixed moments.
@@
 In our case, this is precisely the statement that if then we have that
-which is straightforward.
+which is straightforward. ∎
 We now determine the probability density function for the moments arising in Theorem 1.2. To this end, we recall the definition of the Meijer -function (see Appendix of [EMOT54]).
@@
 where Therefore, we have that
-and our claim follows immediately.
+and our claim follows immediately. ∎
 Finally, we compute the probability density function associated to the moments that appear in Theorem 1.2.
@@
 Computing the integral using (2) of Lemma 7.3, we obtain
-has multiplicative reduction at whereas has good reduction at Therefore, Proposition 5.1 applied to combined with Lemma 5.7 gives the claimed moments.
-This follows from Lemma 2.2 and Proposition 5.5.
+Applying (1) of Lemma 7.3 with and we obtain
+The residue theorem then implies that
+By Lemma 7.4, our claim follows. ∎
+8. Proofs of the Main Results
+Proof of Theorem 1.2.
+Proof of Theorem 1.5.
+Proof of Theorem 1.8.